# latexfontFont shape

### v1

@@ 4.2. ExpTime-hardness of NFA explorability
We will perform here an encoding in the converse direction: starting from an instance $(\mathcal{B},f)$ of the PCP, we build polynomially an NFA $\mathcal{A}$ such that $\mathcal{A}$ is explorable if and only if $(\mathcal{B},f)$ is a negative instance of the PCP.
@@
Let $\mathcal{C}$ be the $4$-state automaton of Section 2.4, that is non-explorable and accepts all words on alphabet $\Sigma_{\mathcal{C}}=\{a,b\}$. Recall that, as an instance of the PCP, $\mathcal{B}$ does not come with an acceptance condition. We will define its accepting set as $F_{\mathcal{B}}=Q_{\mathcal{B}}\setminus\{f\}$.
We will take for $\mathcal{A}$ the product automaton ×$\mathcal{B}\times\mathcal{C}$ on alphabet $\Sigma_{A}=\Sigma_{\mathcal{B}}\times\Sigma_{\mathcal{C}}$, with the union acceptance condition: $\mathcal{A}$ accepts whenever one of its components accepts. The transitions of $\mathcal{A}$ are defined as expected: $(p,p^{\prime})\overset{a_{1},a_{2}}{\longrightarrow}(q,q^{\prime})$ in $\mathcal{A}$ whenever $p\overset{a_{1}}{\longrightarrow}q$ in $\mathcal{B}$ and $p^{\prime}\overset{a_{2}}{\longrightarrow}q^{\prime}$ in $\mathcal{C}$.
@@
Let us give some intuition on why we can obtain a polynomial reduction in one direction, but did not manage to do so in the other direction. Intuitively, the explorability problem is “more difficult” than the PCP for the following reason. In the PCP, Spoiler is allowed to play any letters, and the winning condition just depends on the current position. On the contrary, the winning condition of the $k$-explorability game mentions that the word chosen by Spoiler must belong to the language of the NFA. In order to verify this, we a priori need to append to the arena an exponential deterministic automaton for this language, and this is what is done in Section 4.1. This complicated winning condition is also the source of difficulty in the problem of recognizing HD parity automata.
@@ 4.3. EXPTIME algorithm for [0,2]-explorability
The algorithm was adapted from the ExpTime algorithm for the PCP from [BDG+19]. We will use a variant of this first algorithm, and thus recall here the main ideas of the latter algorithm, and describe how we adapt it to our setting.
Let $\mathcal{A}$ be an NFA, together with a target state $f$. The idea in [BDG+19] is to abstract the population game with arbitrary many tokens by a game called the capacity game. This game allows Determiniser to describe only the support of his set of tokens, i.e. the set of states occupied by tokens. The sequence of states obtained in a play can be analysed via a notion of bounded capacity, in order to detect whether it actually corresponds to a play with finitely many tokens. This notion can be approximated by the more relaxed finite capacity, which is a regular property that is equivalent to bounded capacity in a context where games are finite-memory determined. This property of finite capacity can be verified by a deterministic parity automaton, yielding a parity game that can be won by Spoiler if and only if $(\mathcal{A},f)$ is a positive instance of the PCP. Since this parity game has size exponential in $\mathcal{A}$, this yields an ExpTime algorithm for the PCP.
@@
Our aim is to build a parity game $G$ of size exponential in $\mathcal{A}$ such that deciding whether $\mathcal{A}$ is explorable amounts to deciding the winner of $G$. We want $G$ to be an abstraction of the explorability game, where the number of tokens is not explicitly present. First, we need to control that the infinite word played by Spoiler is in $L(\mathcal{A})$. This requires to build a deterministic parity automaton $\mathcal{D}$ recognizing $L(\mathcal{A})$, and incorporate it into the arena. The size of $\mathcal{D}$ is exponential with respect to $\mathcal{A}$, and the number of priorities is polynomial. We then follow [BDG+19] and build the capacity game augmented with $\mathcal{D}$. In this game, Spoiler plays a letter at each step, and Determiniser chooses a subgraph of the run-DAG of $\mathcal{A}$, that intuitively describes all the transitions taken by his tokens. The condition inherited from the capacity game is enforced as well, in order to force Determiniser to play a run-DAG compatible with a finite number of tokens. Such a DAG is $2$-winning if there are infinitely many steps containing a $2$-transitions. If it is not $2$-winning, it may be unable to have a given token avoid seeing $1$-transitions infinitely often, which means that no run through the support would be accepting. Therefore, Determiniser also has, at all time, a ”challenger” that they want to prevent from seeing a $1$-transition. If they fail at such a task infinitely often, the run is said to be $1$-losing.
By combining all these ingredients, we show that we can as in [BDG+19] obtain a parity game of exponential size characterizing explorability of $\mathcal{A}$, yielding the wanted ExpTime algorithm.
@@
We consider a non-deterministic $[0,2]$-automaton $\mathcal{A}=(\Sigma,Q,q_{0}^{\mathcal{A}},\Delta_{\mathcal{A}},\mathrm{rk})$, where $\mathrm{rk}$ is a function $Q\to[0,2]$ giving the parity rank of each transition.
Let $\mathcal{D}$ be a deterministic parity automaton for $L(\mathcal{A})$, that we can obtain via any standard ExpTime algorithm. We will also make use of the capacity game from [BDG+19], in particular let $\mathcal{T}$ be the deterministic parity automaton built in [BDG+19, Thm 4.5], that accepts a run-DAG if and only if infinitely many tokens are needed to instantiate it. [^3] The notation $\mathcal{T}$ stands for “tracking list automaton” as it is called in [BDG+19]. The alphabet of this automaton consists in transfer graphs of $\mathcal{A}$, i.e. subsets of $Q\times Q$. Both $\mathcal{D}$ and $\mathcal{T}$ have an exponential size and a polynomial number of priorities with respect to the size of $\mathcal{A}$.
@@
We say that Spoiler wins the play if the run $q_{0}^{\mathcal{D}}q_{1}^{\mathcal{D}}q_{2}^{\mathcal{D}}\dots$ of $\mathcal{D}$ is parity accepting, while only finitely many $G_{i}$ contain 2-transitions (from $\Delta_{\mathcal{A}}$) and there is an infinite number of eliminations. Spoiler also wins if the run $q_{0}^{\mathcal{T}}q_{1}^{\mathcal{T}}q_{2}^{\mathcal{T}}\dots$ of $\mathcal{T}$ is parity accepting, witnessing that the sequence $G_{1}G_{2}\dots$ of transfer graphs cannot be instantiated with finitely many tokens.
@@
Proof 4.1.
The winning condition is a boolean condition of parity conditions, hence the game is $\omega$-regular, and thus finite-memory determined.
@@
If Determiniser has a finite-memory winning strategy $\tau$ in the $[0,2]$-capacity game, then he has a strategy $\sigma$ in the $k$-tokens explorability game for some $k$, such that any play consistent with $\sigma$ has its projection consistent with $\tau$. Additionnally, it is possible to chosse $\sigma$ such that the challenger token in the $[0,2]$-capacity game is always instantiated by a particular token in the $k$-explorability game, that can change at each elimination event in the $[0,2]$-capacity game.
Proof 4.2.
The strategy $\tau$ is in particular winning for the original capacity game, as Spoiler wins any play where $\mathcal{T}$ accepts. We can therefore directly import the results from [BDG+19]: the finite memory $m$ of $\tau$ ensures that the capacity of any winning play is actually bounded by a constant depending on $m$. From there, [BDG+19, Lem 3.5] ensures that there is a strategy that moves $k$ tokens to realize the transfer graphs of $\tau$, for some $k$ exponential in $m$. This yields overall a doubly exponential upper bound on the number of tokens needed to instantiate the strategy $\tau$. Let us now describe how to build $\sigma$ in order to additionally ensure that the challenger token is always instantiated by a particular token in the $k$-explorability game, that can only change at an elimination event. Whenever the challenger is instantiated (at the beginning or at each elimination), $\sigma$ will simply choose the token of minimal index $r$ among those in the new challenger state. Then, as all tokens play the same role in the instantiation strategy from [BDG+19], it is always possible to have token $r$ follow the challenger path in $\sigma$, until the next elimination event or forever if there is no more elimination. Indeed, we are guaranteed that the wanted transition is always available, as it is part of the current transfer graph $G$.
@@
Proof 4.4.
Here we do not need any finite-memory determinacy result. Let $\sigma$ be a winning strategy for Determiniser in the $k$-explorability game. We will show that this strategy can be lifted to a winning strategy $\tau$ in the $[0,2]$-capacity game. When playing in the $[0,2]$-capacity game, Determiniser will keep in memory a corresponding play of the $k$-explorability game, and follow the strategy $\sigma$ in it. Then, to answer to the letters played by Spoiler, Determiniser will do the followingc:
@@
  • for the challenger, simply loop over all tokens: when the current challenger is token $i$ and sees a $1$-transition, an elimination event is witnessed, and the challenger is switched to token $i+1$, looping back to 00 after $k-1$.
@@
To conclude and obtain 19, we have to show that the $[0,2]$-capacity game can actually be solved with the wanted complexity.
@@
  • Whenever a $3$-transition is seen on the second component, produce a rejecting $12j+$ 1 rank,
@@ 5.1. Definition and basic results
[$\kappa$-explorability game] Consider an automaton $\mathcal{A}$ and a cardinal $\kappa$. The $\kappa$-explorability game on $\mathcal{A}$ is played on the arena $(Q_{\mathcal{A}})^{\kappa}$, between Determiniser and Spoiler. They play as follows.
@@
We will say in particular that $\mathcal{A}$ is $\omega$-explorable if Determiniser wins the game with $\omega$ tokens. We use here the notation $\omega$ for convenience, it should be understood as the countably infinite cardinal $\aleph_{0}$. We will however explicitly use the fact that such an amount of tokens can be labelled by $\mathbb{N}$, in order to describe strategies for Spoiler or Determiniser in the $\omega$-explorability game. The following lemma gives a first few results on generalized explorability.
@@
We now want to prove that there are automata that are $\omega$-explorable but not explorable. One such automaton is given in Figure 5 (left), where the rejecting sink state is omitted. Against any finite number of tokens, Spoiler has a strategy to eliminate them one by one, by playing $a$ while Determiniser sends tokens to $q_{1}$, and $b$ the first time $q_{1}$ is empty after the play of Determiniser. On the other hand, with tokens indexed by $\omega$, Determiniser can keep the token 00 in $q_{0}$, and send token $i$ to $q_{1}$ at step $i$. Those strategies are winning, which proves both non explorability and $\omega$-explorability of the automaton.
The last item is proven by the second example from Figure 5. A winning strategy for Spoiler against $\omega$ tokens consists in labelling the tokens with integers, then targeting each token one by one (first token 00, then $1$, $2$, etc.). Each token is removed using the correct two-letters sequence ($a$, then $b$ if the token is in $q_{1}$ or $a$ if it is in $q_{2}$). With this strategy, every token is removed at some point, even if there might always be tokens in the game.
$q_{0}$$q_{1}$$q_{2}$$a$$a$$b$$a,b$
$q_{0}$$q_{1}$$q_{2}$$a$$a$$a$$b$
@@ 5.2. ExpTime algorithm for coBüchi automata
To prove this result, we will use the following elimination game. $\mathcal{A}$ will from here on correspond to a coBüchi (complete) automaton. We start by building a deterministic coBüchi automaton $\mathcal{D}$ for $L(\mathcal{A})$ (e.g. using the breakpoint construction [MH84]). We will assume here that the coBüchi condition of $\mathcal{A}$ is state-based to simplify a bit the presentation, but as remarked in the introduction it is straightforward to accomodate transition-based acceptance as well. {defi}[Elimination game] The elimination game is played on the arena $\mathcal{P}(Q_{\mathcal{A}})\times Q_{\mathcal{A}}\times Q_{\mathcal{D}}$. The two players are named Protector and Eliminator, and the game proceeds as follows, starting in the position ($(\{q_{0}^{\mathcal{A}}\},q_{0}^{\mathcal{A}},q_{0}^{\mathcal{D}})$.
@@
To prove this result, we simply need to note that the winning condition is a parity condition of fixed index. If we label the coBüchi states $p_{i}$ of $\mathcal{D}$ with rank $3$, the coBüchi states $q_{i}$ of $\mathcal{A}$ with rank $2$, and the others with $1$, then take the highest rank in $(B_{i},q_{i},p_{i})$ (ignoring $B_{i}$), Eliminator wins if and only if the highest rank appearing infinitely often is $2$. As any parity game with $3$ ranks can be solved in polynomial time [CJK+17], this is enough to get the result. Since parity games are positionnally determined [EJ91], the elimination game is as well.
@@
  • At first, the memory holds the initial state ($(\{q_{0}^{\mathcal{A}}\},q_{0}^{\mathcal{A}},q_{0}^{\mathcal{D}},0)$, and the current challenger is given by the last component via $f$: it is the token $f(0)$.
@@
With Lemmas 29 and 30 we get a proof of Theorem 28, since the elimination game associated to $\mathcal{A}$ is of exponential size and can be built using exponential time.

### v2

@@ 4.2. ExpTime-hardness of NFA explorability
Proof 4.2.
We will perform here an encoding in the converse direction: starting from an instance $(\mathcal{B},f)$ of the population control problem, we build polynomially an NFA $\mathcal{A}$ such that $\mathcal{A}$ is explorable if and only if $(\mathcal{B},f)$ is a negative instance of the population control problem.
@@
Let $\mathcal{C}$ be the $4$-state automaton of Section 2.4, that is non-explorable and accepts all words on alphabet $\Sigma_{\mathcal{C}}=\{a,b\}$. Recall that, as an instance of the population control problem, $\mathcal{B}$ does not come with an acceptance condition. We will define its accepting set as $F_{\mathcal{B}}=Q_{\mathcal{B}}\setminus\{f\}$.
We will take for $\mathcal{A}$ the product automaton $\mathcal{B}\times\mathcal{C}$ on alphabet $\Sigma_{A}=\Sigma_{\mathcal{B}}\times\Sigma_{\mathcal{C}}$, with the union acceptance condition: $\mathcal{A}$ accepts whenever one of its components accepts. The transitions of $\mathcal{A}$ are defined as expected: $(p,p^{\prime})\overset{a_{1},a_{2}}{\longrightarrow}(q,q^{\prime})$ in $\mathcal{A}$ whenever $p\overset{a_{1}}{\longrightarrow}q$ in $\mathcal{B}$ and $p^{\prime}\overset{a_{2}}{\longrightarrow}q^{\prime}$ in $\mathcal{C}$.
@@
Let us give some intuition on why we can obtain a polynomial reduction in one direction, but did not manage to do so in the other direction. Intuitively, the explorability problem is “more difficult” than the population control problem for the following reason. In the population control problem, Spoiler is allowed to play any letters, and the winning condition just depends on the current position. On the contrary, the winning condition of the $k$-explorability game mentions that the word chosen by Spoiler must belong to the language of the NFA. In order to verify this, we a priori need to append to the arena an exponential deterministic automaton for this language, and this is what is done in Section 4.1. This complicated winning condition is also the source of difficulty in the problem of recognizing HD parity automata.
@@ 4.3. EXPTIME algorithm for [0,2]-explorability
The algorithm was adapted from the ExpTime algorithm for the population control problem from [BDG+19]. We will use a variant of this first algorithm, and thus recall here the main ideas of the latter algorithm, and describe how we adapt it to our setting.
Let $\mathcal{A}$ be an NFA, together with a distinguished state $f$. The idea in [BDG+19] is to abstract the population game with arbitrary many tokens by a game called the capacity game. This game allows Determiniser to describe only the support of his set of tokens, i.e. the set of states occupied by tokens. Edges that are taken by at least one token are also specified at each step. Thus, it is possible that such a description is not actually realizable with finitely many tokens, e.g. if a state has infinitely many outgoing edges but only finitely many incoming edges throughout the play. This is captured by a notion of bounded capacity for plays in this game: this notion characterizes whether such a play, given as a sequence of supports and egdes, can be instantiated with finitely many tokens. However, this notion of bounded capacity is not $\omega$-regular. This is why [BDG+19] introduces the more relaxed notion of finite capacity, which is $\omega$-regular, and coincides with bounded capacity in a context of finite-memory strategies.
Thus this property of finite capacity can be verified by a deterministic parity automaton that can be incorporated into the arena. Overall, the capacity game is a parity game that can be won by Spoiler if and only if $(\mathcal{A},f)$ is a positive instance of the population control problem. Since this parity game has size exponential in $\mathcal{A}$, this yields an ExpTime algorithm for the population control problem.
@@
Our aim is to build a parity game $G$ of size exponential in $\mathcal{A}$ such that deciding whether $\mathcal{A}$ is explorable amounts to deciding the winner of $G$. We want $G$ to be an abstraction of the explorability game, where the number of tokens is not explicitly present. First, we need to control that the infinite word played by Spoiler is in $L(\mathcal{A})$. This requires to build a deterministic parity automaton $\mathcal{D}$ recognizing $L(\mathcal{A})$, and incorporate it into the arena. The size of $\mathcal{D}$ is exponential with respect to $\mathcal{A}$, and the number of priorities is polynomial. We then follow [BDG+19] and build the capacity game augmented with $\mathcal{D}$. In this game, Spoiler plays a letter at each step, and Determiniser chooses a subgraph of the run-DAG of $\mathcal{A}$, that intuitively describes all the transitions taken by his tokens. It is then possible to check in an $\omega$-regular way that the run-DAG chosen by Determiniser can actually be instantiated with finitely many tokens. Thus, this game is a faithful abstraction of the $k$-explorability game for arbitrarily high $k$, if we ignore for now the fact that one token has to follow an accepting run. Then, in order to incorporate this condition as well, we define such a DAG to be $2$-winning if there are infinitely many steps containing a $2$-transition, this is a sufficient condition for one of the token to follow an accepting run. If the DAG is not $2$-winning, we have to additionally check that one token through this run-DAG eventually stops seeing $1$-transitions. Therefore, we add the following feature to the game: Determiniser has, at all time, a ”challenger” that is not supposed to see a $1$-transition. If this constraint is failed infinitely often, the run is said to be $1$-losing. Overall, Determiniser has to build a run-DAG that is either $2$-winning, or witness via his challenger token that one run through the DAG is not $1$-losing. This allows us to faithfully abstract the $k$-explorability on $[0,2]$-automata for arbitrarily high $k$, via a fixed finite game of size exponential in $\mathcal{A}$ incorporating all these ingredients. This is how we obtain the ExpTime algorithm for explorability of $[0,2]$-automata.
@@
We consider a non-deterministic $[0,2]$-automaton $\mathcal{A}=(\Sigma,Q,q_{0}^{\mathcal{A}},\Delta_{\mathcal{A}},\mathrm{rk})$, where $\mathrm{rk}$ is a function $\Delta_{A}\to[0,2]$ giving the parity rank of each transition.
Let $\mathcal{D}$ be a deterministic parity automaton for $L(\mathcal{A})$, that we can obtain via any standard ExpTime algorithm. We will also make use of the capacity game from [BDG+19], in particular let $\mathcal{T}$ be the deterministic parity automaton from [^3] This automaton is only implicitly described in [BDG+19] in terms of “tracking lists”, hence our name $\mathcal{T}$ [BDG+19, Thm 4.5], that accepts a run-DAG if and only if it has infinite capacity. The alphabet of this automaton consists in transfer graphs of $\mathcal{A}$, i.e. subsets of $Q\times Q$. Both $\mathcal{D}$ and $\mathcal{T}$ have an exponential size and a polynomial number of priorities with respect to the size of $\mathcal{A}$.
@@
We say that Spoiler wins the play if all the following conditions are simultaneously enforced:
  • the run $q_{0}^{\mathcal{D}}q_{1}^{\mathcal{D}}q_{2}^{\mathcal{D}}\dots$ of $\mathcal{D}$ is parity accepting
  • only finitely many $G_{i}$ contain 2-transitions (from $\Delta_{\mathcal{A}}$)
  • there is an infinite number of eliminations.
Alternatively, Spoiler also wins if the run $q_{0}^{\mathcal{T}}q_{1}^{\mathcal{T}}q_{2}^{\mathcal{T}}\dots$ of $\mathcal{T}$ is parity accepting, witnessing that the sequence $G_{1}G_{2}\dots$ of transfer graphs cannot be instantiated with finitely many tokens.
@@
Proof 4.3.
The winning condition is a Boolean combination of parity conditions, hence the game is $\omega$-regular, and thus finite-memory determined.
@@
If Determiniser has a finite-memory winning strategy $\tau$ in the $[0,2]$-capacity game, then he has a strategy $\sigma$ in the $k$-tokens explorability game for some $k$, such that any play consistent with $\sigma$ has its projection consistent with $\tau$. Additionnally, it is possible to choose $\sigma$ such that the challenger token in the $[0,2]$-capacity game is always instantiated by a particular token in the $k$-explorability game, that can change at each elimination event in the $[0,2]$-capacity game.
Proof 4.4.
The strategy $\tau$ guarantees in particular a finite capacity, as Spoiler wins any play where $\mathcal{T}$ accepts. We can therefore directly import the results from [BDG+19]: the finite memory $m$ of $\tau$ ensures that the capacity of any winning play is actually bounded by a constant depending on $m$. From there, [BDG+19, Lem 3.5] ensures that there is a strategy that moves $k$ tokens to realize the transfer graphs of $\tau$, for some $k$ exponential in $m$. This yields overall a doubly exponential upper bound on the number of tokens needed to instantiate the strategy $\tau$. Let us now describe how to build $\sigma$ in order to additionally ensure that the challenger token is always instantiated by a particular token in the $k$-explorability game, that can only change at an elimination event. Whenever the challenger is instantiated (at the beginning or at each elimination), $\sigma$ will simply choose the token of minimal index $r$ among those in the new challenger state. Then, as all tokens play the same role in the instantiation strategy from [BDG+19], it is always possible to have token $r$ follow the challenger path in $\sigma$, until the next elimination event or forever if there is no more elimination. Indeed, we are guaranteed that the wanted transition is always available, as it is part of the current transfer graph $G$.
@@
Proof 4.6.
Here we do not need any finite-memory determinacy result. Let $\sigma$ be a winning strategy for Determiniser in the $k$-explorability game. We will show that this strategy can be lifted to a winning strategy $\tau$ in the $[0,2]$-capacity game. When playing in the $[0,2]$-capacity game, Determiniser will keep in memory a corresponding play of the $k$-explorability game, and follow the strategy $\sigma$ in it. Then, to answer to the letters played by Spoiler, Determiniser will do the following:
@@
  • for the challenger, simply loop over all tokens: when the current challenger is token $i$ and sees a $1$-transition, an elimination event is witnessed, and the challenger is switched to token $i+1$, looping back to 0 after k1k-1.
@@
To conclude and obtain 16, we have to show that the $[0,2]$-capacity game can actually be solved with the wanted complexity.
@@
  • Whenever a $3$-transition is seen on the second component, produce a rejecting $2j+1$ rank,
@@ 5.1. Definition and basic results
$\kappa$-explorability game] Consider an automaton $\mathcal{A}$ and a cardinal $\kappa$. The $\kappa$-explorability game on $\mathcal{A}$ is played on the arena $(Q_{\mathcal{A}})^{\kappa}$, between Determiniser and Spoiler. They play as follows.
@@
We will say in particular that $\mathcal{A}$ is $\omega$-explorable if Determiniser wins the game with $\omega$ tokens. We use here the notation $\omega$ for convenience, it should be understood as the countably infinite cardinal $\aleph_{0}$. We will however explicitly use the fact that such an amount of tokens can be labelled by $\mathbb{N}$, in order to describe strategies for Spoiler or Determiniser in the $\omega$-explorability game. The following lemma gives a first few results on generalised explorability.
@@
We now want to prove that there are automata that are $\omega$-explorable but not explorable. One such safety automaton is given in Figure 4 (left), where the rejecting sink state is omitted. Against any finite number of tokens, Spoiler has a strategy to eliminate them one by one, by playing $a$ while Determiniser sends tokens to $q_{1}$, and $b$ the first time $q_{1}$ is empty after the play of Determiniser. On the other hand, with tokens indexed by $\omega$, Determiniser can keep the token 0 in $q_{0}$, and send token $i$ to $q_{1}$ at step $i$. Those strategies are winning, which proves both non explorability and $\omega$-explorability of the automaton.
The last item is proven by the second example from Figure 4. A winning strategy for Spoiler against $\omega$ tokens consists in labelling the tokens with integers, then targeting each token one by one (first token 0, then $1$, 22, etc.). Each token is removed using the correct two-letters sequence ($a$, then $b$ if the token is in $q_{1}$ or $a$ if it is in $q_{2}$). With this strategy, every token is removed at some point, even if there might always be tokens in the game.
$q_{0}$$q_{1}$$q_{2}$aaaabba,ba,b
$q_{0}$$q_{1}$$q_{2}$aaaaaabb
@@ 5.2. ExpTime algorithm for coBüchi automata
To prove this result, we will use the following elimination game. $\mathcal{A}$ will from here on correspond to a coBüchi (complete) automaton. We start by building a deterministic coBüchi automaton $\mathcal{D}$ for $L(\mathcal{A})$ (e.g. using the breakpoint construction [MH84]). We will assume here that the coBüchi condition of $\mathcal{A}$ is state-based to simplify a bit the presentation, but as remarked in the introduction it is straightforward to accomodate transition-based acceptance as well. {defi}[Elimination game] The elimination game is played on the arena $\mathcal{P}(Q_{\mathcal{A}})\times Q_{\mathcal{A}}\times Q_{\mathcal{D}}$. The two players are named Protector and Eliminator, and the game proceeds as follows, starting in the position $(\{q_{0}^{\mathcal{A}}\},q_{0}^{\mathcal{A}},q_{0}^{\mathcal{D}})$.
@@
To prove this result, we simply need to note that the winning condition is a parity condition of fixed index. If we label the coBüchi states $p_{i}$ of $\mathcal{D}$ with rank $3$, the coBüchi states $q_{i}$ of $\mathcal{A}$ with rank $2$, and the others with $1$, then take the highest rank in $(B_{i},q_{i},p_{i})$ (ignoring $B_{i}$), Eliminator wins if and only if the highest rank appearing infinitely often is $2$. As any parity game with $3$ ranks can be solved in polynomial time [Zie98], this is enough to get the result. Since parity games are positionnally determined [EJ91], the elimination game is as well.
@@
  • At first, the memory holds the initial state $(\{q_{0}^{\mathcal{A}}\},q_{0}^{\mathcal{A}},q_{0}^{\mathcal{D}},0)$, and the current challenger is given by the last component via $f$: it is the token $f(0)$.
@@
With Lemmas 26 and 27 we get a proof of Theorem 25, since the elimination game associated to $\mathcal{A}$ is of exponential size and can be built using exponential time.